OpenStudy (samigupta8):

three non zero numbers a,b,c are in ap.Increasing a by 1 o increasing c by 2 the number becomes in gp then b equals to

OpenStudy (raden):

what means o there ?

OpenStudy (samigupta8):

it's or

OpenStudy (raden):

ohhh --"

OpenStudy (raden):

a, b, c --> AP b-a = c-b or b = (a+c)/2 ... (1) a+1, b, c ---> GP b/(a+1) = c/b or b^2 = c(a+1) ... (2) a, b, (c+2) ---> GP b/a = (c+2)/b or b^2 = a(c+2) ... (3) (2) = (3), let's see what u get then .......

OpenStudy (samigupta8):

here a =1/2

OpenStudy (samigupta8):

it's not cuming

OpenStudy (raden):

how you get that ? just guessing ?

OpenStudy (samigupta8):

wait it's cuming as c=2a

OpenStudy (raden):

yup. that's right

OpenStudy (samigupta8):

and we get it as b=3/2 a

OpenStudy (raden):

correct again

OpenStudy (samigupta8):

then what

OpenStudy (raden):

b = 3/2 a ---> a = 2b/3 , agree ?

OpenStudy (raden):

c = 2a because a = 2b/3, then the c can rewriten as c = 2a = 2(2b/3) = 4b/3

OpenStudy (raden):

now you have a and c in b

OpenStudy (samigupta8):

ryt

OpenStudy (raden):

subtitute them to one of (2) or (3) equation

OpenStudy (raden):

to (1) equation works too

OpenStudy (raden):

which one do you want :)

OpenStudy (samigupta8):

that's not cuming

OpenStudy (raden):

hmm... yeah the 1st equation doesnt work try use the 2nd equ : b^2 = c(a+1) now subtitute the value of a = (2b)/3 and c = 4b/3

OpenStudy (raden):

got it ?

OpenStudy (samigupta8):

no

OpenStudy (raden):

b^2 = c(a+1) b^2 = (4b)/3 * (2b/3 +1) simplify then solve for b

OpenStudy (samigupta8):

got it

OpenStudy (raden):

really, what is it

OpenStudy (samigupta8):

12

OpenStudy (raden):

smart :)

OpenStudy (samigupta8):

one more question 2 u

OpenStudy (raden):

why to me --"

OpenStudy (samigupta8):

bcuz u hve that intelligence to solve it